What is the probability that Fatima gets fewer heads than tails if she flips 10 coins?
If 10 coins are tossed, you could get 4 heads and 6 tails, 3 heads and 7 tails, 2 heads and 1 tail, 0 heads and 10 tails all giving fewer heads than tails.
Using the binomial distribution ,
P(4 heads) = 10C4 (.5)^4 (.5)^6 = 0.205078.
P(3 heads) = 10C3 (.5)^3 (.5)^7 = 0.117188
P(2 heads) = 10C2 (.5)^2 (.5)^8 = 0.043945
P(1 heads) = 10C1 (.5)^1 (.5)^9 = 0.009766
P(0 heads) =(.5)^10 = 0.000977
Adding all of these probabilities, we have P(fewer heads than tails)= 0.376953
Is probability written in percentage?
It can be, if desired. Probability is a number between 0 and 1. If you want to express that as a percent, simply say between 0% and 100% instead.
How do you calculate distinguishable permutations?
The number of permutations of n distinct objects is n! = 1*2*3* ... *n.
If a set contains n objects, but k of them are identical (non-distinguishable), then the number of distinct permutations is n!/k!.
If the n objects contains j of them of one type, k of another, then there are n!/(j!*k!).
The above pattern can be extended. For example, to calculate the number of distinct permutations of the letters of "statistics":
Total number of letters: 10
Number of s: 3
Number of t: 3
Number of i: 2
So the answer is 10!/(3!*3!*2!) = 50400
What discrete probability distribution can be approximated by a normal dsitribution?
If X has any discrete probability distribution then the sum of a number of observations for X will be normal.
What is the mean mode median standard deviation and standard error of 10 24 35 44 10 and 35?
Mean: 26.33
Median: 29.5
Mode: 10, 35
Standard Deviation: 14.1515
Standard Error: 5.7773
The mean is 0 and the variance is 1. This need not be the case in any other Normal (Gaussian) distribution.
How many continuous probability distributions?
There are an indeterminate number: thousands upon thousands.
What is events for which the outcome of one event does not affect the probability of the other?
Independent events.
How does sample size affect the margin of error?
The larger the sample size, the smaller the margin of error.
This is because the normal distribution has a domain that extends to infinity in both directions.
What is the probability of toss two heads with 2 pennies?
There are four different ways that two coins can land:
T T
T H
H T
H H.
Only one of them is two heads, so if the coins are fair, then
the probability is 1/4 = 25% .
What is the probability of drawing an ace or a king?
There are 52 cards in a deck
there are 4 aces and 4 kings which makes a total of 8 kings and aces.
Assuming that the deck is full and shuffled the probability of drawing an aces or a king is 8/52 which simplifies to 2/13
How many different ways can you arrange 20 photos in an array?
Assuming that the arrays need to be rectangular, the answer is 14,597,412,049,059,800,000 or approx 15 quintillion.
z = (75 - 85)/5 = -10/5 = -2
What is mean by cardinality of sample space?
no of possibilities
for example tossind a fair coin then
the cardinality of sample space is 2
It is 0.0606
The standard z-score such that 80 percent of the distribution is below to the left of this value is?
z = 0.8416
Why is it impossible for the probability to be an improper fraction?
Because the probability of an impossible event is 0 and probability lies within the area of {0-1} meaning 0 is impossible and 1 is definite. If you took an improper fraction such as 5/4 and made it a proper faction, it would be 1 1/4 making it more than 1 or (100%) and that wouldn't make any sense because 4/4 (1) is definite, anything more isn't a probability.
How many eyes are there on the back of a deck of cards?
There need not be any. Most decks of cards have backs that are geometric patterns, not eyes.
What are the similarities of a theoretical and empirical probabilities?
They are both measures of the likelihood of specified events.
Why is t score equal to z score in a normal distribution?
Because as the sample size increases the Student's t-distribution approaches the standard normal.
What is the probability of you getting a 4 of diamonds?
The probability of pulling a four of diamonds from a standard deck of playing cards with the jokers removed is 1:52, the same as pulling any other individual card from the deck.